OpenStudy (anonymous):

The ratio of the angle measures of the acute angles in a right triangle is 1:2

OpenStudy (anonymous):

The ratio of the angle measures of the acute angles in a right angle is 1:2 Let x = the multiplier, the 3 angles always = 180, therefore: 90 + x + 2x = 180 3x = 180 - 90 x = 90/3 x = 30 is the multiplier, therefore the two angles are 30 and 60 degrees !